The problem is: Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
x + y = 4
2x – y = 2
I have no clue how to even begin these types of problems. Any help is greatly appreciated.

Answer by stanbon(75887)

x + y = 4
2x – y = 2
-----------
In this case you can just add the two equations and the y-terms will disappear:
Adding you get:
3x=6
x=2
-------------
Substitute that into one of the original equations to solve for y:
2+y=4
y=2
-----------
Solution:
x=2,y=2
=============
Cheers,
Stan H.